# Olga Zatsarynna1, Emad Bahrami1, Yazan Abu Farha2, Gianpiero Francesca 3, Juergen Gall1, 4

### v1

@@ Abstract
Our work addresses the problem of stochastic long-term dense anticipation. The goal of this task is to predict actions and their durations several minutes into the future based on provided video observations. Anticipation over extended horizons introduces high uncertainty, as a single observation can lead to multiple plausible future outcomes. To address this uncertainty, stochastic models are designed to predict several potential future action sequences. Recent work has further proposed to incorporate uncertainty modelling for observed frames by simultaneously predicting per-frame past and future actions in a unified manner. While such joint modelling of actions is beneficial, it requires long-range temporal capabilities to connect events across distant past and future time points. However, the previous work struggles to achieve such a long-range understanding due to its limited and/or sparse receptive field. To alleviate this issue, we propose a novel MANTA (MAmba for ANTicipation) network. Our model enables effective long-term temporal modelling even for very long sequences while maintaining linear complexity in sequence length. We demonstrate that our approach achieves state-of-the-art results on three datasets—Breakfast, 50Salads, and Assembly101—while also significantly improving computational and memory efficiency.
@@ 1 Introduction
In our work, we focus on stochastic anticipation, where multiple predictions for a single video observation are made to account for the uncertainty of future forecasting. Early works in this area [7, 60] predicted future action sequences based on a fixed set of pre-classified or ground-truth observed actions. Recently, GTDA [55] highlighted the limitations of this approach and proposed to model framewise past and future actions jointly with a single diffusion model, conditioned on the observed frames padded with zeros in place of future entries. To account for the differences between the observed and masked parts of the sequence, they introduced a GTAN diffusion generator - a multi-stage architecture of dilated temporal convolutional layers with data-dependent temporal gates.
Figure 1: We propose a novel generator MANTA that allows for more efficient and effective stochastic long-term anticipation compared to the previous work.
Although their model demonstrated strong performance on the task of stochastic long-term anticipation, the proposed GTAN generator has certain limitations. Firstly, in the early layers of each stage, convolutional kernels have local support due to the small dilation factor. However, given the input structure, where observed frames are followed by masking values, the limited receptive field prevents padded values from accessing observed context and thus updating future frame values. While the multi-stage architecture partially mitigates this, it introduces additional parameters and increases computation. Moreover, while in the later layers, the dilation rate increases the receptive field, it remains sparse due to the small kernel size, with each sequence element accessing only a limited set of other elements, making updates highly restrictive. These limitations are particularly pronounced for longer videos with extended anticipation horizons, where the future-padded part of the input sequence becomes challenging to process.
@@ 3 Method
where $g$ is a feature extractor, $x_{i}$ is the $i^{\text{th}}$ video frame, $\mathcal{Y}^{i}$ is the variable corresponding to frame $i$, and $\tilde{\mathcal{Y}}^{i}$ is a sample from $\mathcal{Y}^{i}$; $P$ and $F$ represent the number of past and future frames, respectively. Recently, Zatsarynna et al. [55] proposed to jointly model the uncertainty of actions in both observed and future frames using a common diffusion model. In this way, they modelled a more general distribution - the joint observation-conditioned probability of both observed and future actions. That is:
$\tilde{\mathcal{Y}}^{1:P+F}$ $\sim P\big{(}\mathcal{Y}^{1:P+F}\mid g(x_{1}):g(x_{P})\big{)}.$ (2)
This allowed their diffusion to model actions in the future and past frames in a consistent manner. Furthermore, to take into account distinctions between these two types of variables, they proposed a GTAN diffusion generator network, that employed gated dilated temporal convolutional layers to process the sequence with data-dependent gates. The use of dilated convolutions for temporal processing, however, poses several limitations. Specifically, the exponentially increasing dilation rates ($2^{l-1}$) of the convolutional kernels result in limited receptive fields in the early layers and sparse receptive fields lacking local context in the later layers. This leads to two issues: the limited receptive field of early layers prevents observed information from reaching future masked entries, while the large, sparse receptive fields of the later layers restrict updates and lack local awareness. Although the multi-stage design intends to address these issues, it only partially mitigates them, while adding to computational and memory costs.
In this work, we address the limitations mentioned above. While we follow Zatsarynna et al. [55] in harnessing the unified diffusion process for observed and future action modelling, we introduce a novel MANTA generator model. By utilizing the Mamba [13]-based layers, our model sustains a global receptive field across the entire network while preserving local awareness. Additionally, relying on the data-dependent gating our model processes the input sequence adaptively, considering the difference between its two components - observed and future actions. In the following sections, we present the details of our proposed model. In Sec. 3.1, we describe how the diffusion process is formulated for addressing the task of stochastic long-term dense anticipation. Then, in Sec. 3.2, we present the details of our proposed MANTA model.
@@ 3.1 Diffusion for Long-Term Dense Anticipation
General Diffusion. Diffusion models are generative models that learn to produce samples from data distributions by reconstructing progressively noisier versions of data points. To achieve this, two Markov chain processes are defined: the forward and the reverse. The forward Markov process $q(\mathcal{Y}_{t}|\mathcal{Y}_{t-1})$ defines the transition from the data distribution to Gaussian noise, using a predefined noising schedule. In contrast, the reverse Markov process $p_{\theta}(\mathcal{Y}_{t-1}|\mathcal{Y}_{t})$ models the inverse transition, from random noise back to the data distribution.
The reverse process is parameterized by a diffusion generator network $\mathcal{G}_{\theta}$ that is trained to reconstruct data points corrupted by the forward process: ${\tilde{\mathcal{Y}}_{0,t}}=\mathcal{G}_{\theta}(\mathcal{Y}_{t},t)$. Here, $\mathcal{Y}_{t}$ is the corrupted data sample and $\tilde{\mathcal{Y}}_{0,t}$ is the reconstructed data point.
Anticipation Diffusion. We follow the formulation of diffusion anticipation model proposed in [55]. Following Eq. (2), anticipation diffusion reconstructs one-hot encoded per-frame past and future ground-truth actions conditioned on the padded visual features of observed frames. More specifically, the diffusion latent variables $\mathcal{Y}_{t}$ are introduced to model per-frame actions, with a total of $(P+F)$ variables. To condition anticipation on past video frames, latent variables $\mathcal{Y}_{t}$ are concatenated with a vector $\mathcal{X}$, constructed by extending observed visual features with zeros in place of future action variables to account for the corresponding unobservable visual features:
@@
Figure 2: (Top) Overview of the proposed MANTA model. (Bottom) Illustration of the structure of the (a) MANTA block, (b) Bidirectional State-Space Layer, as well as (c) traversal directions of the BSSL for the temporal input sequence.
@@ 3.2 Model
The overview of our proposed MANTA model is shown in Fig. 2. The key contribution of our model is the Bidirectional Selective State-Space Layer (BSSL), which enables effective and efficient long-range temporal modelling and supports selective sequence processing through data-dependent gating. These properties are essential for long-term dense anticipation, where framewise prediction of past and future actions requires reasoning over lengthy sequences with a structured pattern: observed frames followed by masked future frames. Next, we describe our proposed BSS layer in Sec. 3.2.2, after detailing deep state-space models in Sec. 3.2.1, which serve as its foundation. We then discuss the overall structure of our proposed MANTA model in Sec. 3.2.3.
@@ 3.2.1 Deep Structured State-Space Models
where $\textbf{A}\in\mathbb{R}^{N\times N}$, $\textbf{B}\in\mathbb{R}^{N\times 1}$ and $\textbf{C}\in\mathbb{R}^{N\times 1}$ are state, input and output matrices, respectively. Deep SSMs are defined by transforming the above formulation into a discrete-time learnable system. To this end, continuous-time equations (4) and (5) are discretized with a time-step parameter $\Delta$.
@@
These discretized equations form the foundation for a class of deep SSMs S4 [14], which define a sequence-to-sequence transformation parametrized by a set of learnable parameters (A, B, C, and $\Delta$) following Eq. (6) and (7). To compute this model, its parameters are transformed into ($\bar{\textbf{A}}$, $\bar{\textbf{B}}$) using Eq. (8) and (9), after which output $y$ is acquired either via linear recurrence or global convolution.
While S4 models maintain a global receptive field, they do not implement selective value processing, since the parameters (A, B, C, and $\Delta$) remain fixed over time. Recently, Gu et al. [13] proposed an improved S6 layer by introducing the selection mechanism into the S4 model. To this end, they proposed to make parameters input-dependent by computing them as functions of $x\in\mathbb{R}^{B\times T\times D}$, i.e. $\textbf{B}(x)\in\mathbb{R}^{B\times T\times N}$, $\textbf{C}(x)\in\mathbb{R}^{B\times T\times N}$, and $\Delta(x)\in\mathbb{R}^{B\times T\times D}$. This increases the model’s adaptability, enabling data-driven information selection from both hidden states $h_{t}$ and input sequence entries $x_{t}$. While it also leads to increased capacity, the output of S6 model can be efficiently computed using a parallel-scan algorithm.
The tradeoff between keeping and pruning the information in the input and hidden vectors is controlled by the parameters of S6, which can be interpreted to have specific functions within the model, as described in [13, 18]. On one hand, $\Delta$ can be viewed as an input gate, controlling how much information from the input $x_{t}$ gets propagated into the hidden state $h_{t}$. Conversely, $\bar{\textbf{A}}$ functions as a forget gate, determining the extent of pruning of the previous hidden state $h_{t-1}$. Finally, C controls how much information from $h_{t}$ propagates into the output $y_{t}$. In this way, the data-dependent gating of the S6 adopts a similar role as the gated convolution of the GTAN, which adjusts the output of temporal convolutional layers using temporal gates.
@@ 3.2.2 Bidirectional State-Space Layer
For BSSL, we make use of the above-described S6 layer as our temporal processing unit, due to its attractive properties of maintaining locally-aware [18] global receptive field, in contrast to GTAN, whose receptive field is limited by the dilation factor of the convolution layers. Apart from that, the selective modelling of the input sequence based on the data-dependent gating of the underlying SSM model is a crucial property which enables MANTA to perform anticipation effectively.
We base our layer on the Mamba block, proposed in  [13], which incorporates the S6 layer into the residual block for efficient and effective sequence processing. The general overview of the BSSL’s structure is depicted in Fig. 2(b). Following [13], our proposed layer consists of two branches: the main branch (top) and the residual branch (bottom). The residual branch implements a multiplicative residual connection within the layer. Specifically, it uses a linear projection layer to map the input feature $F_{t}^{b}$ to a space with $2$-times expanded channels, followed by the activation function $\sigma$ (SiLU [19, 38]), resulting in the vector $R_{t}^{b}$:
$R_{t}^{b}=\sigma(FF(F_{t}^{b}))$ (10)
@@
As shown in Fig. 2(b), the forward and backward paths are applied to the linearly projected input sequence separately. Both paths follow the same flow: after a depth-wise convolution, the input is processed by the S6 layer, whose output is then activated by $\sigma$. After that, a residual connection is applied to the outputs of both paths independently. Importantly, before and after processing the sequence with the backward path, it is flipped in time. Formally:
@@
Here $H^{K}$ is a 1D convolutional filter with size $K$, flipping across temporal dimension is denoted as $\leftarrow$, while * and $\odot$ are convolution and element-wise multiplication, respectively. The output feature of the BSSL block is computed by accumulating the outputs of backward and forward paths as follows:
@@ 3.2.3 MANTA
Depicted in Fig. 2 (top), our proposed network consists of $B$ MANTA blocks that process the input to predict future actions. Given pre-extracted observed frame features, we first pad them with zeros in place of future frames as described in Eq. (3), creating a conditioning vector $\mathcal{X}\in\mathbb{R}^{(P+F)\times n_{d}}$. $\mathcal{X}$ is then concatenated with latent variables $\mathcal{Y}_{t}$ along the channel dimension to form $F_{t}\in\mathbb{R}^{(P+F)\times(n_{c}+n_{d})}$, which is sequentially processed by MANTA blocks to generate the output feature $\tilde{F}_{t}$, based on which the final prediction $\tilde{\mathcal{Y}}_{0,t}\in\mathbb{R}^{(P+F)\times n_{c}}$ is made using an MLP layer. Here $n_{c}$ and $n_{d}$ denote number of classes and number of feature channels, respectively.
@@ 4.1 Datasets
To evaluate our proposed model, we use three datasets commonly utilized for stochastic long-term anticipation tasks: Breakfast [24], Assembly101 [42] and 50Salads [50]. Breakfast [24] is a dataset of cooking videos in which human actors prepare 10 different breakfast meals in various kitchen environments. It comprises 1,712 videos that are densely annotated with temporal segments across 48 action classes. The dataset includes videos of varying durations, with the longest sequence lasting 10.8 minutes, which requires an anticipation horizon of up to 5.4 minutes. Following previous work, we evaluate our model using 4 predefined cross-validation splits and report the mean performance across these splits.
@@
4.2 Implementation Details
Following previous methods, we make use of the same pre-extracted frame features. Specifically, we use I3D features from [1, 11] for the Breakfast and 50Salads datasets, while for Assembly101 we rely on TSM-features [28] provided by [42]. Additionally, to ensure a fair comparison with [55], we closely follow their implementation. We provide further details in the supp. mat.

### v2

# Olga Zatsarynna1,4, Emad Bahrami1,4, Yazan Abu Farha2, Gianpiero Francesca3, Juergen Gall1,4

@@ Abstract
Long-term dense action anticipation is very challenging since it requires predicting actions and their durations several minutes into the future based on provided video observations. To model the uncertainty of future outcomes, stochastic models predict several potential future action sequences for the same observation. Recent work has further proposed to incorporate uncertainty modelling for observed frames by simultaneously predicting per-frame past and future actions in a unified manner. While such joint modelling of actions is beneficial, it requires long-range temporal capabilities to connect events across distant past and future time points. However, the previous work struggles to achieve such a long-range understanding due to its limited and/or sparse receptive field. To alleviate this issue, we propose a novel MANTA (MAmba for ANTicipation) network. Our model enables effective long-term temporal modelling even for very long sequences while maintaining linear complexity in sequence length. We demonstrate that our approach achieves state-of-the-art results on three datasets—Breakfast, 50Salads, and Assembly101—while also significantly improving computational and memory efficiency. Our code is available at https://github.com/olga-zats/DIFF_MANTA.
@@ 1 Introduction
In our work, we focus on stochastic anticipation, where multiple predictions for a single video observation are made to account for the uncertainty of future forecasting. Early works in this area [7, 61] predicted future action sequences based on a fixed set of pre-classified or ground-truth observed actions. Recently, GTDA [57] highlighted the limitations of this approach and proposed to model framewise past and future actions jointly with a single diffusion model, conditioned on the observed frames padded with zeros in place of future entries. To account for the differences between the observed and masked parts of the sequence, GTDA introduced a GTAN diffusion generator - a multi-stage architecture of dilated temporal convolutional layers with data-dependent temporal gates.
Figure 1: We propose a novel MANTA diffusion generator that allows for more efficient and effective stochastic long-term anticipation compared to the previous work.
Although it achieves strong performance on the task of stochastic long-term anticipation, the proposed GTAN generator has certain limitations. Firstly, in the early layers of each stage, convolutional kernels have local support due to the small dilation factor. However, given the input structure, where observed frames are followed by masking values as depicted in Fig. 1, the limited receptive field prevents padded values from accessing observed context and thus updating future frame values. While the multi-stage architecture partially mitigates this, it introduces additional parameters and increases computation. Moreover, while in the later layers, the dilation rate increases the receptive field, it remains sparse due to the small kernel size, with each sequence element accessing only a limited set of other elements, making updates highly restrictive. These limitations are particularly pronounced for longer videos with extended anticipation horizons, where the future-padded part of the input sequence becomes challenging to process.
@@ 3 Method
where $g$ is a feature extractor, $x_{i}$ is the $i^{\text{th}}$ video frame, $\mathcal{Y}^{i}$ is the variable corresponding to frame $i$, and $\tilde{\mathcal{Y}}^{i}$ is a sample from $\mathcal{Y}^{i}$; $P$ and $F$ represent the number of past and future frames, respectively. In [57], the uncertainty of actions in both observed and future frames is jointly modelled by a more general distribution - the joint observation-conditioned probability of both observed and future actions:
$\tilde{\mathcal{Y}}^{1:P+F}$ $\sim P\big{(}\mathcal{Y}^{1:P+F}\mid g(x_{1}):g(x_{P})\big{)},$ (2)
and a common diffusion model predicts actions in the future and past frames in a consistent manner. Furthermore, to take into account distinctions between these two types of variables, [57] proposed a GTAN diffusion generator network, that employs dilated gated temporal convolutional layers to process the sequence with data-dependent gates. The use of dilated convolutions for temporal processing, however, poses several limitations. Specifically, the exponentially increasing dilation rates ($2^{l-1}$) of the convolutional kernels result in limited receptive fields in the early layers and sparse receptive fields lacking local context in the later layers. This leads to two issues: the limited receptive field of early layers prevents observed information from reaching future masked entries, while the large, sparse receptive fields of the later layers restrict updates and lack local awareness. Although the multi-stage design intends to address these issues, it only partially mitigates them, while adding to computational and memory costs.
In this work, we address the limitations mentioned above. While we follow Zatsarynna et al. [57] in harnessing the unified diffusion process for observed and future action modelling, we introduce a novel MANTA generator model. By utilizing Mamba [13]-based layers, our model sustains a global receptive field across the entire network while preserving local awareness. Additionally, relying on the inherent data-dependent gating, our model processes the input sequence adaptively, considering the difference between its two components - observed and future actions. In the following sections, we present the details of our proposed model. In Sec. 3.1, we describe how the diffusion process is formulated for addressing the task of stochastic long-term dense anticipation. Then, in Sec. 3.2, we present the details of our proposed MANTA model.
@@ 3.1 Diffusion for Long-Term Dense Anticipation
General Diffusion. Diffusion models are generative models that learn to produce samples from data distributions by reconstructing progressively noisier versions of data points. To achieve this, two Markov chain processes are defined: the forward and the reverse. The forward Markov process $q(\mathcal{Y}_{t}|\mathcal{Y}_{t-1})$ defines the transition from the data distribution to Gaussian noise, using a predefined noising schedule. In contrast, the reverse Markov process $p_{\theta}(\mathcal{Y}_{t-1}|\mathcal{Y}_{t})$ models the inverse transition, from random noise back to the data distribution. The reverse process is parameterized by a diffusion generator network $\mathcal{G}_{\theta}$ that is trained to reconstruct data points corrupted by the forward process: ${\tilde{\mathcal{Y}}_{0,t}}=\mathcal{G}_{\theta}(\mathcal{Y}_{t},t)$. Here, $\mathcal{Y}_{t}$ is the corrupted data sample and $\tilde{\mathcal{Y}}_{0,t}$ is the reconstructed data point.
Anticipation Diffusion. We adopt the formulation of the diffusion anticipation model proposed in [57]. Following Eq. (2), anticipation diffusion reconstructs one-hot encoded per-frame past and future ground-truth actions conditioned on the padded visual features of observed frames. More specifically, the diffusion latent variables $\mathcal{Y}_{t}$ are introduced to model per-frame actions, with a total of $(P+F)$ variables. To condition anticipation on past video frames, latent variables $\mathcal{Y}_{t}$ are concatenated with a vector $\mathcal{X}$, constructed by extending observed visual features with zeros in place of future action variables to account for the corresponding unobservable visual features:
@@
Figure 2: (Top) Overview of the proposed MANTA model. Given a noise vector $\mathcal{Y}_{t}$ and a conditioning vector $\mathcal{X}$, constructed by extending the features of the $P$ observed frames with zero padding in place of the $F$ future frames, we concatenate and forward them through our proposed MANTA model. As output, MANTA predicts action classes for both observed and future frames. (Bottom) Illustration of the structure of the (a) MANTA block, (b) Bidirectional State-Space Layer, as well as (c) traversal directions of the BSSL for the temporal input sequence.
@@ 3.2 Model
The overview of our proposed MANTA model is shown in Fig. 2. The key component of our model is the Bidirectional Selective State-Space Layer (BSSL), which enables effective and efficient long-range temporal modelling and supports selective sequence processing through data-dependent gating. These properties are essential for long-term dense anticipation, where framewise prediction of past and future actions requires reasoning over long sequences with a structured pattern: observed frames followed by masked future frames. Next, we describe our proposed BSS layer in Sec. 3.2.2, after detailing deep state-space models in Sec. 3.2.1, which serve as its foundation. We then discuss the overall structure of our proposed MANTA model in Sec. 3.2.3.
@@ 3.2.1 Deep Structured State-Space Models
where $\textbf{A}\in\mathbb{R}^{N\times 1}$, $\textbf{B}\in\mathbb{R}^{N\times 1}$ and $\textbf{C}\in\mathbb{R}^{N\times 1}$ are state, input and output matrices, respectively. Deep SSMs are defined by transforming the above formulation into a discrete-time learnable system. To this end, continuous-time equations (4) and (5) are discretized with a time-step parameter $\Delta$ in the following way:
@@
These discretized equations form the foundation for a deep SSM called S4 [15], which defines a sequence-to-sequence transformation parametrized by (A, B, C, and $\Delta$) following Eq. (6) and (7). To compute the output of this model, its parameters are transformed into ($\bar{\textbf{A}}$, $\bar{\textbf{B}}$) using Eq. (8) and (9), after which output $y$ is acquired either via linear recurrence or global convolution.
While S4 models maintain a global receptive field, they do not implement selective value processing, since the parameters (A, B, C, and $\Delta$) remain fixed over time. Recently, Gu et al. [13] proposed an improved S6 layer by introducing the selection mechanism into the S4 model. To this end, they proposed to make parameters input-dependent by computing them as functions of $x\in\mathbb{R}^{B\times T\times D}$, i.e., $\textbf{B}(x)\in\mathbb{R}^{B\times T\times N}$, $\textbf{C}(x)\in\mathbb{R}^{B\times T\times N}$, and $\Delta(x)\in\mathbb{R}^{B\times T\times D}$. This increases the model’s adaptability, enabling data-driven information selection from both hidden states $h_{t}$ and input sequence entries $x_{t}$.
This selection is controlled by the parameters of S6, which can be interpreted to have specific functions within the model, as described in [13, 18]. Parameter $\Delta$ can be viewed as an input gate, controlling how much information from the input $x_{t}$ gets propagated into the hidden state $h_{t}$. The parameter $\bar{\textbf{B}}$ similarly influences the degree of contribution of $x_{t}$ to $h_{t}$. Conversely, $\bar{\textbf{A}}$ functions as a forget gate, determining the extent of pruning of the previous hidden state $h_{t-1}$. Finally, C controls how much information from $h_{t}$ propagates into the output $y_{t}$.
In this way, the data-dependent gating of the S6 adopts a similar role as the gated convolution of the GTAN, which adjusts the output of temporal convolutional layers using gating. While data-dependent processing also leads to increased capacity of S6, its output can be efficiently computed using a parallel-scan algorithm.
@@ 3.2.2 Bidirectional State-Space Layer
For BSSL, we make use of the above-described S6 layer as our temporal processing unit, due to its attractive properties of maintaining a locally-aware [18] global receptive field, in contrast to GTAN, whose receptive field is limited by the dilation factor of its convolution layers. Apart from that, the selective modelling of the input sequence based on the data-dependent gating of the underlying SSM model is crucial for enabling MANTA to perform anticipation effectively.
We base our layer on the Mamba block, proposed in  [13], which incorporates the S6 layer into the residual block for efficient and effective sequence processing. The general overview of BSSL’s structure is depicted in Fig. 2(b). Following [13], our proposed layer consists of two branches: the main branch (top) and the residual branch (bottom). The residual branch implements a multiplicative residual connection within the layer. Specifically, it uses a linear projection layer to map the input feature $F_{t}^{b}$ to a space with $2$-times expanded channels, followed by the activation function $\sigma$ (SiLU [19, 39]), resulting in the vector $R_{t}^{b}$:
$R_{t}^{b}=\sigma(FF(F_{t}^{b})).\vspace{-0.2cm}$ (10)
@@
As shown in Fig. 2(b), the forward and backward paths are applied to the linearly projected input sequence separately. Both paths follow the same flow: after a depth-wise convolution, its output is activated using $\sigma$ and is then processed by the S6 layer. After that, a residual connection is applied to the outputs of both paths independently. Importantly, before and after processing the sequence with the backward path, it is flipped in time. Formally:
@@
Here $H^{K}$ is a 1D convolutional filter with size $K=3$, flipping across temporal dimension is denoted as $\leftarrow$, while * and $\odot$ are convolution and element-wise multiplication, respectively. The output feature of the BSSL block is computed by accumulating the outputs of backward and forward paths as follows:
@@ 3.2.3 MANTA
Depicted in Fig. 2 (top), our proposed network consists of $B$ MANTA blocks that process the input to predict future actions. Given pre-extracted observed frame features, we first pad them with zeros in place of future frames as described in Eq. (3), creating a conditioning vector $\mathcal{X}\in\mathbb{R}^{(P+F)\times n_{d}}$. $\mathcal{X}$ is then concatenated with latent variables $\mathcal{Y}_{t}$ along the channel dimension to form $F_{t}\in\mathbb{R}^{(P+F)\times(n_{c}+n_{d})}$, which is sequentially processed by MANTA blocks to generate the output feature $\tilde{F}_{t}$, based on which the final prediction $\tilde{\mathcal{Y}}_{0,t}\in\mathbb{R}^{(P+F)\times n_{c}}$ is made using an MLP layer. Here, $n_{c}$ and $n_{d}$ denote number of classes and number of feature channels, respectively.
@@ 4.1 Datasets
To evaluate our proposed model, we use three datasets, commonly utilized for stochastic long-term anticipation tasks: Breakfast [25], Assembly101 [44] and 50Salads [51].
Breakfast [25] is a dataset of cooking videos in which human actors prepare 10 different breakfast meals in various kitchen environments. It comprises 1,712 videos that are densely annotated with temporal segments across 48 action classes. The dataset includes videos of varying durations, with the longest sequence lasting 10.8 minutes, which requires an anticipation horizon of up to 5.4 minutes. Following previous work, we evaluate our model using 4 predefined cross-validation splits and report the mean performance across these splits.
@@
Following previous methods, we make use of the same pre-extracted frame features. Specifically, we use I3D features from [2, 11] for the Breakfast and 50Salads datasets, while for Assembly101 we rely on TSM features [29] provided by [44]. Additionally, to ensure a fair comparison with [57], we closely follow their implementation. We provide further details in the supp. mat.